# Gaze-guided Hand-Object Interaction Synthesis: Dataset and Method

### v4

@@ Introduction
Since the hand diffusion effectively captures the overall distribution of hand-object interactions, it falls short of achieving precision in finer details, such as maintaining contact or avoiding penetrations. Therefore, we introduce HOI-Manifold Guidance to strengthen the physics constraint while preserving the original data distribution. We define three guidance functions from the perspective of Kinematic, Contact and Penetration.
(1) Hand-Object Kinematic Guidance.
@@
(2) Hand-Object Contact Guidance.
@@
(3) Hand-Object Penetration Guidance.
@@
$\mathcal{L}_{\text{penetration}}=-\sum_{i=1}^{l}\text{min}(0,\tilde{\mathbf{F}}^{i}\cdot\mathbf{n}^{i}+\eta),$ (8)
@@
The naive guidance strategy, which directly adds the computed gradient $\nabla_{{\mathcal{H}}^{t}}\mathcal{L}(\hat{\mathcal{H}}^{0})$ to the sample, may misalign the final motions with the original data distribution, resulting in unnatural motions. Therefore, we aim to develop a guidance strategy that can achieve guidance targets while preserving the original distribution. Recently, a conditional diffusion model with the Spherical Gaussian constraint, named DSG (Yang et al. 2024) is proposed to offer larger, adaptive step sizes while preserving the original data distribution. However, its effectiveness has mainly been shown in image-generation tasks. Motivated by DSG, we introduce the Spherical Gaussian constraint to our sampling process, which ensures that the generated hand motion remains within the clean data distribution, resulting in natural and smooth hand motion generation.
@@
where $w$ represents the guidance rate that lies in [0,1] and $\mathcal{D}$ represents the weighted direction. After sampling, we use post-optimization to get the corresponding MANO parameters $\hat{\mathbf{H}}$. Details are illustrated in the Appendix.
Motion Selection via Gaze-Interaction Consistency Score
Motions generated by diffusion do not always fully conform to the gaze distribution. To address this, we explore gaze-interaction consistency from both a local and a global perspective, as shown in Figure 2(b). Locally, inspired by Desanghere and Marotta (2015), we observe that areas receiving more visual attention are more likely to be contacted. So we compute the contact map $\mathbf{M}_{c}=e^{-\alpha\cdot d(\mathbf{P}_{h},\mathbf{P}_{o})}$ and gaze map $\mathbf{M}_{g}=e^{-\alpha\cdot d(\mathbf{G},\mathbf{P}_{o})}$ separately by the distance from hand vertices $\mathbf{P}_{h}$ and gaze point $\mathbf{G}$ to the object surface $\mathbf{P}_{o}$. The gaze-contact consistency score $S_{c}$ is then calculated as follows,
@@
where $\mathbf{G}_{\delta}$ is the set of object points index that indicates where the distance $d(\mathbf{G},\mathbf{P}_{o})<\delta$, and $|\mathbf{G}_{\delta}|$ is the number of elements in this set. Globally, gaze-hand-object trajectories are expected to exhibit inherent relationships. To quantify this, we use Dynamic Time Warping (DTW) (Seshan 2022) to measure the gaze-hand-object trajectory similarity,
@@
Experiments
@@
Table 2: Quantitative comparison between baselines and our method. * indicates an adapted version of the original method tailored to our task. - refers to the same results as Ours, as OMOMO uses the same Stage-1 module to acquire object motions. The ground truth Diversity is $4.227$ and diversity results are better if the metric is closer to the real distribution.
@@
Data Split and Evaluation Metrics
Comparison
Our method is the first to generate both hand and object motions from gaze data, and there are no direct methods available for comparison. The most closely related state-of-the-art work is Text2HOI (Cha et al. 2024), which generates a contact map conditioned on texts and then combines it to produce hand-object motions. We found that the gaze map and contact map are consistent, allowing us to adapt this approach by generating the contact map conditioned on gaze. We also employed MDM (Tevet et al. 2022) as a baseline, as it is a classic benchmark for diffusion-based motion generation. To better extract gaze features and for a fairer comparison, both MDM and Text2HOI use our designed spatial-temporal gaze encoding. Additionally, Object-Conditioned Hand Motion Synthesis is a key component of our method. Although there is no direct baseline for our entire task, OMOMO (Li, Wu, and Liu 2023) is closely related to this aspect, so we compare our approach with OMOMO, using our fully designed stage-1 to acquire the object motions. Quantitative and qualitative results are shown in Table 2 and Figure 3, demonstrating that our method significantly outperforms the other baselines.
Ablation Study
Results on Seen and Unseen Objects.
As shown in Figure 3, our method outperforms others on both seen and unseen objects, delivering motions that are natural and plausible, while closely aligning with the underlying information encoded in the gaze data. A more detailed quantitative comparison will be provided in the Appendix.
Gaze Representation and Encoding.
@@
As the first method to use gaze as a condition for motion generation, we explore the effectiveness of different gaze representations and spatial-temporal feature encoding. The results are in Table 3. The results show that our method effectively captures both spatial and temporal features, enabling efficient extraction of the rich information encoded in gaze data.
@@
Table 3: Ablation on Gaze Representation and Encoding. Ours- means our full method without motion selection for a fair comparison.
HOI-Manifold Guidance.
The guidance-free method suffers from problems like lack of contact and penetration, while the naïve guidance strategy partially mitigates these problems but introduces motion distortions, compromising the naturalness of the results. In contrast, our HOI-Manifold Guidance strategy strikes a balance between these challenges, achieving superior outcomes as shown in Table 4.
@@
Table 4: Ablation on Different Guidance Strategies. Ours- means our full method without motion selection for a fair comparison.
Motion Selection via Gaze-interaction Consistency Score.
Table 5 demonstrates the effectiveness of our designed gaze-interaction consistency score, showing that motions selected by this score are more closely aligned with the information implied in the gaze data.
@@
w/o select 117.9 157.6 154.3 4.25
w/o global 117.9 156.4 145.7 4.40
w/o local 117.5 148.9 149.8 4.35
Ours 117.4 146.5 144.8 4.75
Table 5: Ablation on Motion Selection via Gaze-interaction Consistency Score.
Conclusion

### v5

@@ IV-B2 Stage-2: Hand Kinematic Synthesis with HOI Guidance
We define $\mathcal{B}(\cdot)$as a function that calculates the bone lengths, taking the hand joints $\mathbf{J}$ as input. The bone length loss $\mathcal{L}_{\text{bone}}$ is defined as:
$\mathcal{L}_{\text{bone}}=||\mathcal{B}(\mathbf{J})-\mathcal{B}(\hat{\mathbf{J}})||_{2}.$ (9)
Since hand diffusion effectively captures the overall distribution of hand-object interactions, it falls short of achieving precision in finer details, such as maintaining contact or avoiding penetrations. Therefore, we introduce HOI-Manifold Guidance to strengthen the physics constraint while preserving the original data distribution. We define three guidance functions from the perspective of Kinematic, Contact and Penetration.
Hand-Object Kinematic Guidance.
@@
Hand-Object Contact Guidance.
@@
Hand-Object Penetration Guidance.
@@
$\mathcal{L}_{\text{penetration}}=-\sum\nolimits_{i=1}^{l}\text{min}(0,\tilde{\mathbf{F}}^{i}\cdot\mathbf{n}^{i}+\eta),$ (12)
@@
The naive guidance strategy, which directly adds the computed gradient $\nabla_{{\mathcal{H}}^{t}}\mathcal{L}(\hat{\mathcal{H}}^{0})$ to the sample, may misalign the final motions with the original data distribution, resulting in unnatural motions. Therefore, we aim to develop a guidance strategy that can achieve guidance targets while preserving the original distribution. Recently, a conditional diffusion model with the Spherical Gaussian constraint, named DSG [52] is proposed to offer larger, adaptive step sizes while preserving the original data distribution. However, its effectiveness has mainly been shown in image-generation tasks. Motivated by DSG, we introduce the Spherical Gaussian constraint to our sampling process, which ensures that the generated hand motion remains within the clean data distribution, resulting in natural and smooth hand motion generation.
@@
where $w$ represents the guidance rate that lies in [0,1] and $\mathcal{D}$ represents the weighted direction.
Since we represent the hand pose using hand joints, a post-optimization process is required to derive the corresponding MANO representation. We first calculate the global R by matching the actual MCP joint positions of the hand with its ideal position in the static state. The MCP joints are crucial for determining the global orientation of the hand because they serve as the primary pivot points for finger movement and are centrally located, providing a stable reference for the overall hand pose. By matching these joints, we can accurately capture the hand’s global rotation, ensuring that the subsequent pose adjustments maintain anatomical correctness. Then we define a loss function $\mathcal{L}_{\text{joint}}$ as optimization target to align the generated joint $\mathbf{J}$ with the joint calculated by manolayer $\mathcal{M}(\cdot)$,
$\mathcal{L}_{\text{joint}}=||\mathbf{J}-\mathcal{M}(\mathbf{H})||_{2}.$ (17)
To preserve the natural and realistic movement of the hand, we apply pose angle constraints at the end of each iteration. These constraints limit the rotation range of hand joints across different axes, with strict controls specifically on the Y-axis rotation of the MCP joints and the Z-axis rotation of the other joints.
IV-C Motion Selection via Gaze-Interaction Consistency Score
Motions generated by diffusion do not always fully conform to the gaze distribution. To address this, we explore gaze-interaction consistency from both a local and a global perspective, as shown in Figure 2(b). Locally, inspired by [3], we observe that areas receiving more visual attention are more likely to be contacted. So we compute the contact map $\mathbf{M}_{c}=e^{-\alpha\cdot d(\mathbf{P}_{h},\mathbf{P}_{o})}$ and gaze map $\mathbf{M}_{g}=e^{-\alpha\cdot d(\mathbf{G},\mathbf{P}_{o})}$ separately by the distance from hand vertices $\mathbf{P}_{h}$ and gaze point $\mathbf{G}$ to the object surface $\mathbf{P}_{o}$. The gaze-contact consistency score $S_{c}$ is then calculated as follows,
@@ IV-C Motion Selection via Gaze-Interaction Consistency Score
where $\mathbf{G}_{\delta}$ is the set of object points index that indicates where the distance $d(\mathbf{G},\mathbf{P}_{o})<\delta$, and $|\mathbf{G}_{\delta}|$ is the number of elements in this set. Globally, gaze-hand-object trajectories are expected to exhibit inherent relationships. To quantify this, we use Dynamic Time Warping (DTW) [53] to measure the gaze-hand-object trajectory similarity,
@@
V Experiments
V-A Data Split and Evaluation Metrics
V-B Implementation Details
[…]
TABLE III: Quantitative comparison between baselines and our method. The ground truth Diversity is $4.227$ and diversity results are better if the metric is closer to the real distribution.
@@ V-B Implementation Details
V-C Baseline Design and Comparison
Our method is the first to generate both hand and object motions from gaze data, and there are no direct methods available for comparison. Here are some related baselines.
V-C1 Text2HOI*
The most closely related state-of-the-art work is Text2HOI [47], which generates a contact map conditioned on texts and then combines it to produce hand-object motions. Due to the observation of the relationship between gaze and contact, we adapt the approach by replacing the text encoder with our spatial-temporal gaze encoding to generate the contact map. The architecture of the models, as well as the training loss functions, are consistent with those described in the original Text2HOI paper. Since the original authors only released the testing code and did not provide the full training code, we reproduced the training process based on the methodology described in the Text2HOI paper. The definition of the loss function was also referenced from the Contact2Grasp [34] and InterGen [54] code mentioned in Text2HOI.
V-C2 MDM*
We also employed MDM [50] as a baseline, as it is a classic benchmark for diffusion-based motion generation. We replace the condition encoder of MDM [50] with our spatial-temporal gaze encoding and introduce three additional loss functions to better align with our task requirements:
$\mathcal{L}_{\text{MDM*}}=\lambda_{\gamma}\mathcal{L}_{\text{simple}}+\lambda_{o}\mathcal{L}_{\text{obj}}+\lambda_{h}\mathcal{L}_{\text{hand}}+\lambda_{f}\mathcal{L}_{\text{offset}},$ (21)
where $\mathcal{L}_{\text{obj}}$ and $\mathcal{L}_{\text{hand}}$ independently constraint the object and hand translation, while $\mathcal{L}_{\text{offset}}$ is defined to constraint the offset between object center points and hand wrist joints.
$\mathcal{L}_{\text{obj}}=\sum\nolimits_{i=1}^{L}||\mathbf{T}_{o}^{i}-\hat{\mathbf{T}}_{o}^{i}||_{2},$ (22)
$\mathcal{L}_{\text{hand}}=\sum\nolimits_{i=1}^{L}||\mathbf{T}_{h}^{i}-\hat{\mathbf{T}}_{h}^{i}||_{2},$ (23)
$\mathcal{L}_{\text{offset}}=\sum\nolimits_{i=1}^{L}||(\mathbf{T}_{h}^{i}-\mathbf{T}_{o}^{i})-(\hat{\mathbf{T}}_{o}^{i}-\hat{\mathbf{T}}_{h}^{i})||_{2},$ (24)
[…]
V-C3 OMOMO*
Object-Conditioned Hand Motion Synthesis is a key component of our method. Although there is no direct baseline for our entire task, OMOMO [42] is closely related to this aspect, so we compare our approach with OMOMO, using our fully designed stage-1 to acquire the object motions.
Quantitative and qualitative results are shown in Table III and Figure 3, demonstrating that our method significantly outperforms the other baselines across all metrics and visualizations. The MPVPE and FOL metrics for OMOMO* are marked with ”-” because OMOMO* only includes comparisons for stage-2 and directly uses our stage-1 results. Therefore, these two metrics are excluded from the comparison.
TABLE IV: Quantitative comparison on Seen Objects and Unseen Objects.
Methods Seen Objects Unseen Objects
[…]
MDM* 147.5 214.4 165.8 42.30 2.17 4.05 3.55 146.6 231.8 166.8 68.19 2.62 3.40 3.35
Text2HOI* 316.7 324.6 267.0 46.00 1.97 3.00 2.70 327.3 327.8 255.6 63.44 2.66 2.85 2.50
OMOMO* - - 157.2 52.00 1.95 4.30 4.2 - - 163.3 65.57 2.65 4.00 3.60
Ours 113.1 133.3 146.6 59.28 1.92 4.90 4.75 118.7 150.2 144.3 70.55 2.61 4.70 4.55
V-D Ablation Study
V-D1 Results on Seen and Unseen Objects
To evaluate the model’s generalization capabilities and mitigate potential overfitting to object-specific traits, the test set is carefully designed to include a majority of objects not present in the training set. Specifically, among the 275 test sequences, 215 feature entirely new objects unseen during training, while 60 involve objects encountered during training but are paired with more complex interaction tasks, such as assembling or playing chess. These challenging tasks introduce additional errors in hand pose generation, providing a rigorous test of the method’s robustness and adaptability across varying levels of interaction complexity. As shown in Figure 3 and Table IV, our method outperforms others on both seen and unseen objects, delivering motions that are natural and plausible, while closely aligning with the underlying information encoded in the gaze data. The results demonstrate that our method exhibits superior performance on both seen and unseen objects.
V-D2 Gaze Representation and Encoding
@@
As the first method to use gaze as a condition for motion generation, we explore the effectiveness of different gaze representations and spatial-temporal feature encoding. The results are in Table V. The results show that our method effectively captures both spatial and temporal features, enabling efficient extraction of the rich information encoded in gaze data.
TABLE V: Ablation on Gaze Representation and Encoding. Ours- means our full method without motion selection for a fair comparison.
@@
TABLE VI: Ablation on Different Guidance Strategies. Ours- means our full method without motion selection for a fair comparison.
@@
V-D3 HOI-Manifold Guidance
The guidance-free method suffers from problems like lack of contact and penetration, while the naïve guidance strategy partially mitigates these problems but introduces motion distortions, compromising the naturalness of the results. In contrast, our HOI-Manifold Guidance strategy strikes a balance between these challenges, achieving superior outcomes as shown in Table VI.
TABLE VII: Ablation on Motion Selection via Gaze-interaction Consistency Score.
@@ V-D3 HOI-Manifold Guidance
w/o select 117.9 157.6 154.3 65.77 2.51 0.048 4.25 4.45
w/o global 117.9 156.4 145.7 67.89 2.47 0.045 4.40 4.55
w/o local 117.5 148.9 149.8 65.96 2.50 0.046 4.35 4.45
Ours 117.4 146.5 144.8 68.10 2.46 0.044 4.75 4.60
V-D4 Motion Selection via Gaze-interaction Consistency Score
Table VII demonstrates the effectiveness of our designed gaze-interaction consistency score, showing that motions selected by this score are more closely aligned with the information implied in the gaze data.
TABLE VIII: Performance of baseline models on the Hand-Object Motion Reconstruction task within our GazeHOI dataset
[…]
Validation ArcticNet-SF [21] 111.2/43.5 23.7 20.3
ArcticNet-LSTM [21] 114.9/45.2 24.0 22.5
Keypoint Transformer [55] 108.9/42.6 32.6 28.4
Test ArcticNet-SF [21] 65.1/42.6 22.8 29.7
ArcticNet-LSTM [21] 68.9/42.4 23.1 33.0
Keypoint Transformer [55] 111.4/42.5 27.6 44.9
VI Task: Hand-Object Motion Reconstruction
Due to our rigorous data collection process and the precise alignment achieved across multiple coordinate systems, the GazeHOI dataset ensures that the 3D hand-object data is exceptionally well-aligned with the corresponding image data. This high level of alignment makes GazeHOI an ideal resource for tackling complex hand-object reconstruction tasks, as it provides a seamless integration between 3D spatial information and visual cues from images.
VI-A Task Definition
Given a video of hand-object interactions and the object’s mesh, the task goal is to reconstruct continuous, smooth, and natural hand-object interaction motions for each frame.
VI-B Data Split and Metrics
The data utilized in the experiment comprises 12 side views from our complete dataset, totaling 3.0M images. We divide the data by subjects, 8 subjects for training, 1 for evaluation, and 1 for testing. To measure hand-object reconstruction quality, we compute Mean Relative-Root Position Error (MRRPE) for left-to-right and right-to-object interactions and Mean Per Joint Position Error (MPJPE) in local space. Additionally, following [21], we evaluate errors in object poses through a success rate metric that is the percentage of predicted object vertices with an L2 error to the ground truth which is less than 5% of the object’s diameter.
VI-C Baselines and Results
To validate the quality of our dataset, substantiate its utility beyond synthesis tasks for reconstruction tasks, and establish a benchmark for later works, we select two Hand-Object Motion Reconstruction baselines consistent with our dataset setting. The first baseline ArcticNet[21] is a naive model that obtains hand and object motions separately. It has two versions: an image-based model (ArcticNet-SF) and a video-based model (ArcticNet-LSTM) that allows reasoning of hand and object motions. Another one Keypoint Transformer[55] is an image-based model that performs well on several challenging datasets, including InterHand2.6M [56], HO-3D[16] and H2O-3D[55], through solving joint identification. We evaluate its performance on the results of each frame like the ArcticNet-SF. Table VIII shows the quantitative results of our baselines.
VII Conclusion